Hi my name is Liz and I need help to solve this: y=5X-2 I need to solve for X first than for Y, so I can plot at least 3 points on my graph. Thank you
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

y=5X-2 I need to solve for X first than for Y, so I can plot at least 3 points on my graph.
----
Solve for "x":
y = 5x-2
5x = y+2
x = (y+2)/5
-------------
Solve for "y":
y = 5x-2
If x = 0,y = -2 giving you point (0,-2)
If x = 2, y = 5*2-2 ; y = 8 giving you point (2,8)
================================
Plot the two points and draw a line thru them.
graph(400,300,-10,10,-10,10,5x-2)}}
================================
Cheers,
Stan H.
========

Answer by mananth(16946)

Hi my name is Liz and I need help to solve this: y=5X-2 I need to solve for X first than for Y, so I can plot at least 3 points on my graph. Thank you
...
y=5X-2
x=0 y=-2 (0,-2)
y=0,x=2/5 (2/5,0)
x=1,y=3 (1,3)

...
m.ananth@hotmail.ca